OpenStudy (anonymous):

A 250-kg snowmobile is traveling north at 12 m/s. Suppose that 15 s later it is going south at 18 m/s. a. What is the acceleration of the snowmobile? b. What is its momentum at the end of the 15 s?

OpenStudy (anonymous):

Hey There. this is a physics or mechanics question. You should have posted in the afforementioned subjects. a. From what I think, we first find the relative speed, which is the difference in speeds from both directions, being 18-12= 6 m/s. then acceleration is given by speed/time or 6/15=2/5 or 0.4 m/s^2. \ b.Momentum is given by mass*velocity and that could be 250*3= 750Kgms^-1